REU: SCHEDULE FOR THE FIRST TWO WEEKS

There will be an opening meeting Monday, June 17, at 9:00. Pizza and coke will be available around lunchtime that day.

There will be no YSP or SESAME duties the first week. There will be morning sessions during the second week.

Date  YSP/SESAME  Babai  Sally  Farb
Monday, June 17  10:00–12:00, 1:00–2:30 3:00–5:00
Tuesday, June 18 10:00–12:00 1:00–3:00
Wednesday, June 19  10:00–12:00, 1:00–2:30 3:00–5:00
Thursday, June 20 10:00–12:00 1:00 –3:00
Friday, June 21  10:00–12:00, 1:00–2:30 3:00–5:00
Monday, June 24  9:00–12:00  1:00–2:45  3:15–5:00
Tuesday, June 25  9:00–12:00  1:00–3:00
Wednesday, June 26  9:00–12:00  1:00–2:45  3:15–5:00
Thursday, June 27  9:00–12:00  1:00–3:00
Friday, June 28  9:00–12:00  1:00–2:45  3:15–5:00

Graduate student sessions will be organized around this schedule.

Here is the schedule in summary form:

First week:

Opening meeting: 9:00–10:00 M Babai 10:00–12:00 M–F and 1:00–2:30 MWF Farb 1:00– 3:00 TTh Sally 3:00– 5:00 MWF

Second week:

YSP/SESAME 9:00–12:00 M–F Babai 1:00– 2:45 MWF Farb 1:00– 3:00 TTh Sally 3:15– 5:00 MWF